# GOAL: A Generalist Combinatorial Optimization Agent Learner

### v1

@@ Abstract
Machine Learning-based heuristics have recently shown impressive performance in solving a variety of hard combinatorial optimization problems (COPs). However they generally rely on a separate neural model, specialized and trained for each single problem. Any variation of a problem requires adjustment of its model and re-training from scratch. In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving multiple COPs and which can be fine-tuned to solve new COPs. GOAL consists of a single backbone plus light-weight problem-specific adapters, mostly for input and output processing. The backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features. Additionally, problems which involve heterogeneous nodes or edges, such as in multi-partite graphs, are handled through a novel multi-type transformer architecture, where the attention blocks are duplicated to attend only the relevant combination of types while relying on the same shared parameters. We train GOAL on a set of routing, scheduling and classic graph problems and show that it is only slightly inferior to the specialized baselines while being the first multi-task model that solves a variety of COPs. Finally we showcase the strong transfer learning capacity of GOAL by fine-tuning or learning the adapters for new problems, with only few shots and little data.
@@ 1 Introduction
Combinatorial Optimization problems are extensively studied in the Operations Research community, because they are both challenging in terms of theoretical complexity and central in many applications such as transportation, logistics and finance. Traditional methods include exact solvers (typically based on integer programming [32]) and approximation algorithms [55] which provide optimality or approximation guarantees, respectively, but do not scale; and heuristic algorithms [9] which produce good-quality solutions faster but without guarantees. Effective heuristics heavily rely on expert knowledge and are generally problem-specific. A more recent and promising family of methods for CO is based on machine learning, either to learn the hyper-parameters of classic CO solvers, or to directly learn heuristics in an end-to-end fashion [6, 11]. In particular constructive neural heuristics have been successful for solving a variety of CO problems, including routing [5, 42, 30, 33], scheduling [58, 34] and graph problems [27, 45]. However, in most cases, a separate model is specialized and trained for each problem, and sometimes even for each distribution of instances within a problem, especially for varying instance sizes. While much effort has been devoted to improving the generalization to different instance distributions [41, 59, 49, 16], generalization to different problems is less studied. Indeed, although there are strong neural heuristics for many CO problems, any variation of these problems requires adjustment of their model and re-training from scratch. While it is true that some problem-specific model tailoring is always required, since the structure of problem instances and solutions are very diverse, many problems are related and it would be interesting to exploit their commonalities. In this paper, we seek to answer the following question: Can one model be pretrained in order to be easily specialized to solve a variety of CO problems?
The idea of pretraining a unified model on some pretext task so that it can be efficiently fine-tuned to several downstream tasks has been extremely successful in the language and vision domains [15, 10, 46]. In these domains, a self-supervised pretraining task exploits the intrinsic structure and regularities of images and text to learn generic representations. In contrast, in CO, while most problem instances can be modeled as graphs, there is no generic regularity to leverage by a self-supervised task. Graphs, unlike images or text, are a “poor” modality, with little redundancy or intrinsic semantics.
@@
In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving a variety of CO problems and which can be fine-tuned to solve new related problems. GOAL consists of a single backbone and light-weight problem-specific adapters for input and output processing. To handle the large discrepancy between CO problems, we propose two complementary strategies. First, the backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features. Additionally, problems which involve heterogeneous nodes or edges, which typically arise with multipartite graphs, are handled through a novel multi-type transformer architecture, where the backbone mixed-attention blocks are duplicated to attend each relevant combination of types while relying on the same shared parameters. GOAL is able to handle any CO problem whose instances can be represented as a graph and where feasible solutions can be constructed iteratively, as a sequence of node selections. We find that this pattern is generic enough to cover a wide range of problems. We train GOAL in a multi-task setting, by imitation of expert trajectories of eight classic CO problems, spanning multiple domains: Routing problems: the Asymmetric Traveling Salesman Problem (ATSP), the Capacitated Vehicle Routing Problem (CVRP), the CVRP with Time Windows (CVRPTW), the Orienteering Problem (OP); Scheduling problems: the Job Shop Scheduling Problem (JSSP) and the Unrelated Machine Scheduling Problem (UMSP); Packing problems: the Knapsack Problem (KP); and Graph problems: the Minimum Vertex Covering Problem (MVC). The expert trajectories are based on good-quality solutions, easily generated by traditional problem-specific solvers for instances of medium size (generally 100 nodes).
@@
Experimentally we first show that GOAL performs very well on all the training tasks, making it the first model that solves such a variety of CO problems. Second, we illustrate GOAL’s strong zero (or few) shot generalization performance to various distributions shifts from the training tasks’ instances. Third, we fine-tune GOAL on four new problems and demonstrate that it achieves a good performance quickly using little data. We also discuss scenarios where fine-tuning is less effective, hinting at the required conditions for successful transfer; namely that the new task is close enough to at least one of the training tasks to share a similar instance structure, while its objective and constraints can be completely different. We further provide an ablation study to evaluate the effectiveness of our multi-type transformer architecture and mixed-attention blocks. Notably our architecture gives state-of-the-art results on 6 of our 8 training tasks when used in a single-task version of GOAL. In summary, our contributions are the following:
@@
  • We introduce GOAL: a generic and flexible model for end-to-end combinatorial optimization. Two key ingredients of GOAL are the simple yet effective form of mixed-attention blocks which enables handling problems represented by graphs with any combination of node, edge and instance-level features and its multi-type transformer architecture which allows to seamlessly tackle problems with multipartite graph structures.
@@
  • We showcase the potential of such a multi-task pretrained model to be efficiently adapted to handle new CO tasks and provide empirical insights on how closely related the new tasks should be to the training tasks to ensure successful transfer.
@@ 2 Related Works
Our work falls into the category of end-to-end constructive heuristics for combinatorial optimization. Constructive neural heuristics have been successfully applied to a wide range of COPs, including many variants of routing problems [30, 57, 28, 31, 18, 52], scheduling problems [58, 34], as well as classic COPs such as the Knapsack Problem [33] and the Maximum Independent Set [27, 52]. Another successful category of neural CO methods are improvement heuristics, which focus on iteratively enhancing a given complete solution, typically by using a neural network to guide the selection of local search operators [13, 22, 40, 56]. The seminal work of [27] introduces graph embeddings which do work for a variety of CO problems but with limited performance. Current state-of-the-art models are tailored for certain problem structures (e.g. Euclidian routing problems, where nodes are represented by their coordinates plus some extra node or graph features such as demand or vehicle capacity), and successfully handled by transformer-based architectures [30]. For non-euclidian routing problems, [34] and [16] propose adapted architectures to deal with the distance matrix, but they are specialized for such problems and do not apply to those with node-only features. In GOAL, we propose a novel mixed-attention block which can flexibly deal with both node and edge features, incurring no overhead when one or the other is missing.
@@
While the first seminal models for neural CO had a dramatic drop in performance when tested on out-of-training distribution instances [26, 41], several recent approaches show a greatly improved generalization performance, especially to larger instances [18, 35, 49, 43], to distribution shifts of other instance parameters [8], or both [41, 8, 59, 16]. Concurrently with our work, some papers have started to explore the cross-problem generalization for the family of routing problems. Specifically, [36] propose to pretrain the AM model [30] on the TSP and to fine-tune either the entire model or some light-weight modules on the Price Collecting TSP and OP. Additionally, [60] and [37] explore generalization to different variants of the CVRP. They build on the idea of zero-shot compositionality [48] by training a model on CVRPs with some combinations of constraints from a predefined set (e.g. time windows and duration limit constraints) and show that the model generalizes to some unseen combinations of the same constraints. Although these approaches tackle multiple problems, it is important to note that the structures of the considered Euclidian routing problems are very similar and indeed many of those problems were solved by the seminal Attention Model of [30] with little adaptations of the base model. Our paper goes further and considers multi-task training and transfer learning of a new model on a variety of COPs with different structure.
@@
Multi-task (pre-)training has recently shown promising results in related fields. For example, [25] proposed a graph neural network processor capable of learning to execute a variety of algorithms, such as sorting and searching. The paper focuses only on polynomial problems and learns to imitate the execution of classical algorithms. In contrast, we address NP-hard problems and learn by imitation of the solution provided by (potentially exponential) oracle algorithms. Recent approaches in control rely on sequence modeling of expert trajectories from multiple tasks [50, 51]. Although they treat various tasks, with potentially diverse state and action spaces, these works still consider image observations and exploit this structure. [47] proposes GATO, a general-purpose agent obtained by training a large transformer-based model on a hundreds of tasks spanning vision, language and control. It demonstrates the potential of a single network (albeit with 1.2B parameters) to effectively tackle a wide variety of tasks.
@@ 3 Multi-Task Learning for Combinatorial Optimization
Figure 1: Left: Architecture of GOAL for single-type problems; the green components are task-specific, while the others belong to the backbone. Right: Architecture of GOAL for multi-type problems (with here 2 types of nodes T1,T2); “self” and “cross” are identical multi-head mixed attention blocks (same parameters for all types and tasks); “FF” for the two types are also identical Feed Forward blocks; for the edges, all the type combinations are shown (here T11,T12,T21,T22) but each task can define its own. The same set of parameters is used for the two architectures.
@@ 3.1 Constructive CO problem solving
We assume that a CO problem instance is given by a pair of a finite, non-empty set $X$ of feasible solutions, and a real-valued objective function $f$ whose domain contains $X$. Solving the instance consists in finding an optimal solution, i.e. an element of $\arg\min_{x\in X}f(x)$. Constructive approaches assume that $f$ is defined not only over $X$ but over a set of partial solutions, which contains $X$ and is amenable to a step-wise construction procedure: each partial solution can be obtained by a sequence of construction steps. Thus, solving the problem instance can be formulated as a sequential decision procedure which iteratively selects the construction steps so as to reach not only a feasible but an optimal solution. That in turn can be modeled as a Markov Decision Problem (MDP) associated with the problem instance.
The action space of this MDP is naturally defined as the (discrete) set of construction steps, but there are many ways to choose its state space. For the very broad class of so called “tail-recursive” CO problems, [16] shows that an effective choice is the space of all possible instances of the problem. Indeed, the construction of a solution can be modeled as a trajectory in the space of problem instances, following the tail sub-problems [7] obtained after each construction step. All the classical standard NP-hard CO problems, such as vehicle routing problems, machine scheduling, various graph problems etc. are tail-recursive, sometimes modulo a slight reformulation (e.g. TSP is not per se tail-recursive, but path-TSP is). In the sequel, we deal exclusively with tail-recursive CO problems and their construction MDPs, called BQ-MDP in [16].
@@ 3.2 Multi-Task CO Problems
Now, given a class $\mathcal{T}$ of tail-recursive CO problems called tasks, let $\boldsymbol{\Omega}_{t}$ for each task $t{\in}\mathcal{T}$ denote the parametric space used to represent its instances. Consider the new (multi-task) CO problem defined as the union of the tasks in $\mathcal{T}$ (each being a set of instances). Its instances can be represented in the disjoint union $\biguplus_{t\in\mathcal{T}}\boldsymbol{\Omega}_{t}$ consisting of pairs $(t,\omega)$ of a task identifier $t{\in}\mathcal{T}$ and a point $\omega{\in}\boldsymbol{\Omega}_{t}$ in its corresponding parameter space. It is easy to show that the multi-task problem is tail-recursive whenever each of its components is. And its BQ-MDP is then simply the disjoint union of the BQ-MDPs of its components.
In this paper, we study alternatives for the multi-task policy space, beyond the trivial one where no parameter is shared between the policies of the different tasks (Cartesian product of the policy spaces of the individual tasks). The underlying assumption is that if the tasks are sufficiently similar, learning one task may elicit implicit skills useful to other tasks, which are then leveraged by parameter sharing. In the extreme, some tasks may require little training data and/or just a few shot fine-tuning to obtain acceptable performance, leveraging their similarity to more thoroughly trained tasks. Since each task has its own parametric space used to represent its instances, some part of any multi-task policy has to be task specific, and trained by processing at least a few instances of each task: true zero-shot generalization is not possible. The situation is close to multi-lingual Natural Language Processing, where the vocabulary of each language needs to be seen at least once at training. And we borrow the solution used in that context, namely adapters [4], to capture the task-specific parts of the policy.

### v2

# GOAL: A Generalist Combinatorial Optimization Agent Learning

@@ Abstract
Machine Learning-based heuristics have recently shown impressive performance in solving a variety of hard combinatorial optimization problems (COPs). However they generally rely on a separate neural model, specialized and trained for each single problem. Any variation of a problem requires adjustment of its model and re-training from scratch. In this paper, we propose GOAL (for Generalist combinatorial Optimization Agent Learning), a generalist model capable of efficiently solving multiple COPs and which can be fine-tuned to solve new COPs. GOAL consists of a single backbone plus light-weight problem-specific adapters for input and output processing. The backbone is based on a new form of mixed-attention blocks which allows to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features. Additionally, problems which involve heterogeneous types of nodes or edges are handled through a novel multi-type transformer architecture, where the attention blocks are duplicated to attend the meaningful combinations of types while relying on the same shared parameters. We train GOAL on a set of routing, scheduling and classic graph problems and show that it is only slightly inferior to the specialized baselines while being the first multi-task model that solves a wide range of COPs. Finally we showcase the strong transfer learning capacity of GOAL by fine-tuning it on several new problems. Our code is available at https://github.com/naver/goal-co/.
@@ 1 Introduction
Combinatorial Optimization (CO) problems are extensively studied in the Operations Research community, because they are both challenging in terms of theoretical complexity and central in many applications such as transportation, logistics and finance. Traditional CO methods include exact solvers (typically based on integer programming [korte_combinatorial_2018]) and approximation algorithms [vazirani_approximation_2003] which provide optimality or approximation guarantees, respectively, but do not scale; and heuristic algorithms [boussaid_survey_2013] which produce good-quality solutions faster but without guarantees. Effective heuristics heavily rely on expert knowledge and are generally problem-specific. A more recent and promising family of methods for CO is based on Machine Learning, either to learn the hyper-parameters of classic CO solvers, or to directly learn heuristics in an end-to-end fashion [bengio_machine_2021, cappart_combinatorial_2021]. In particular constructive neural heuristics have been successful at solving a variety of CO problems, including routing [bello_neural_2017, nazari_reinforcement_2018, kool_attention_2019, kwon_pomo_2020], scheduling [zhang_learning_2020, kwon_matrix_2022] and graph problems [khalil_learning_2017, qiu_dimes_2022]. However, in most cases, a separate model is specialized and trained for each problem, and sometimes even for each distribution of instances within a problem, especially for varying instance sizes. While much effort has been devoted to improving the generalization to different instance distributions [manchanda_generalization_2022, zhou_towards_2023, son_meta-sage_2023, drakulic_bq-nco_2023], generalization to different problems is less studied. Indeed, although there are strong neural heuristics for many CO problems, any variation of these problems requires adjustment of their model and re-training from scratch. While it is true that some problem-specific model tailoring may be required, since the structure of problem instances and solutions are very diverse, many problems are related and it would be interesting to exploit their commonalities. In this paper, we seek to answer the following question: Can we design and train a single generalist model so that it can be efficiently adapted to solve a variety of CO problems? The idea of pretraining a unified model on some pretext task so that it can be efficiently fine-tuned to several downstream tasks has been extremely successful in the language and vision domains [devlin_bert_2019, brown_language_2020, radford_learning_2021]. In these domains, a self-supervised pretraining task exploits the intrinsic structure and regularities of images and text to learn generic representations. In contrast, in CO, while most problem instances can be modeled as graphs, there is no generic regularity to leverage by a pretext task. Graphs, unlike images or text contain little redundancy or intrinsic semantics.
@@
In this paper, we propose a new model, named GOAL (for Generalist combinatorial Optimization Agent Learning), which consists of a single backbone and light-weight problem-specific adapters for input and output processing. To handle the large discrepancy between CO problems, we propose three complementary strategies. First, to ensure learning related representations for different CO problems, we propose to use a problem-specific projection of the input instance features to a small fixed-size representation, which is then projected into the main embedding space using a shared codebook. Second, to handle problems defined on graphs with arbitrary combinations of node, edge and instance-level features, we design a new form of mixed-attention blocks for the backbone, injecting edge information at the core of each attention block. Finally, we introduce a novel multi-type transformer architecture to deal with problems which involve heterogeneous types of nodes or edges (typically represented as multipartite graphs). The idea is to duplicate the backbone’s mixed-attention blocks in order to attend each meaningful combination of types while relying on the same fixed set of shared parameters. In principle, GOAL is able to handle any CO problem whose instances can be represented as a graph and where feasible solutions can be constructed iteratively, as a sequence of node selections. We find that this pattern is generic enough to cover a wide range of problems. We train GOAL in a multi-task setting, by imitation of expert trajectories of eight classic CO problems, spanning multiple domains: Routing problems: the Asymmetric Traveling Salesman Problem, the Capacitated Vehicle Routing Problem with or without Time Windows, the Orienteering Problem ; Scheduling problems: the Job Shop Scheduling Problem and the Unrelated Machine Scheduling Problem; Packing problems: the Knapsack Problem; and Graph problems: the Minimum Vertex Covering Problem. The expert trajectories are based on good-quality solutions, easily generated by traditional problem-specific solvers for instances of medium size (generally 100 nodes).
@@
Experimentally we first show that GOAL performs very well on all the training tasks, making it the first model that solves such a variety of CO problems. Second, we consider eight new problems and demonstrate that GOAL can be fine-tuned to reach a good performance, either in a few minutes using few-shot imitation learning or in a couple of hours without supervision. In both cases, the fine-tuning takes a fraction of the time and data needed to train a model from scratch. We further provide an ablation study to evaluate the effectiveness of a shared codebook, multi-type transformer architecture and mixed-attention blocks. Notably the GOAL architecture gives state-of-the-art results on 7 out of our 8 training tasks when used in as single-task model. In summary, our contributions are the following:
@@
  • We introduce GOAL: a generic and flexible model for end-to-end combinatorial optimization. Two key ingredients of GOAL are the simple yet effective form of mixed-attention blocks which enables handling problems represented by graphs with any combination of node-, edge- and instance-level features and its multi-type transformer architecture which allows to seamlessly tackle problems with multipartite graph structures.
@@
  • We showcase the potential of such a multi-task pretrained model to be efficiently adapted to handle varied new CO tasks, with or without supervision. We believe this is a key step toward a foundation model for CO.
@@ 2 Related Works
Our work falls into the category of end-to-end constructive heuristics for combinatorial optimization. Constructive neural heuristics have been successfully applied to a wide range of COPs, including many variants of routing problems [kool_attention_2019, xin_multi-decoder_2021, kim_sym-nco_2022, kool_deep_2021, fu_generalize_2021, sun_difusco_2023], scheduling problems [zhang_learning_2020, kwon_matrix_2022], as well as classic COPs such as the Knapsack Problem [kwon_pomo_2020] and the Maximum Independent Set [khalil_learning_2017, sun_difusco_2023]. Another successful category of neural CO methods are improvement heuristics, which focus on iteratively enhancing a given complete solution, typically by using a neural network to guide the selection of local search operators [chen_learning_2019, hottung_neural_2019, ma_learning_2021, wu_learning_2020]. The seminal work of [khalil_learning_2017] introduces graph embeddings which do work for a variety of CO problems but with limited performance. Current state-of-the-art models are tailored for certain problem structures (e.g. euclidian routing problems, where nodes are represented by their coordinates plus some extra node or graph features such as demand or vehicle capacity), and successfully handled by transformer-based architectures [kool_attention_2019]. For non-euclidian routing problems, [kwon_matrix_2022] and [drakulic_bq-nco_2023] propose adapted architectures to deal with the distance matrix, but they are specialized for such problems and do not apply to those with node-only features. In GOAL, we propose a novel mixed-attention modules which can flexibly deal with both node and edge features, incurring no overhead when one or the other is missing.
@@
While the first seminal models for neural CO had a dramatic drop in performance when tested on out-of-training distribution instances [joshi_learning_2022, manchanda_generalization_2022], several recent approaches show a greatly improved generalization performance, especially to larger instances [fu_generalize_2021, li_learning_2021, son_meta-sage_2023, pan_h-tsp_2023], to distribution shifts of other instance parameters [bi_learning_2022], or both [manchanda_generalization_2022, bi_learning_2022, zhou_towards_2023, drakulic_bq-nco_2023]. Concurrently with our work, some papers have started to explore the cross-problem generalization for the family of routing problems. Specifically, [lin_cross-problem_2024] propose to pretrain the AM model [kool_attention_2019] on the TSP and to fine-tune either the entire model or some light-weight modules on the Price Collecting TSP and OP. Additionally, [zhou_mvmoe_2024, liu_multi-task_2024, berto_routefinder_2024] explore generalization to different variants of the CVRP. They build on the idea of zero-shot compositionality [ruis_independent_2021] by training a model on CVRPs with some combinations of constraints from a predefined set (e.g. time windows and duration limit constraints) and show that the model generalizes to some unseen combinations of the same constraints. Although these approaches tackle multiple problems, it is important to note that the considered Euclidian routing problems remain very similar in structure and indeed many of those problems were solved by the seminal Attention Model of [kool_attention_2019] with little adaptations of its base model. Our paper goes further and considers multi-task training and transfer learning of a new model on a variety of COPs with different structures.
@@
Multi-task (pre-)training has recently shown promising results in related fields. For example, [ibarz_generalist_2022] proposed a graph neural network processor capable of learning to execute a variety of algorithms, such as sorting and searching. The paper focuses only on polynomial problems and learns to imitate the execution of classical algorithms. In contrast, we address NP-hard problems and learn by imitation of the solution provided by (potentially exponential) specialized solvers. Recent approaches in control rely on sequence modeling of expert trajectories from multiple tasks [stooke_decoupling_2021, sun_smart_2023]. Although they treat various tasks, with potentially diverse state and action spaces, these works still consider image observations and exploit this common structure. [reed_generalist_2022] proposes GATO, a general-purpose agent obtained by training a large transformer-based model on hundreds of tasks spanning vision, language and control. It demonstrates the potential of a single network (albeit with 1.2B parameters) to effectively tackle a wide variety of tasks.
@@ 3 Multi-Task Learning for Combinatorial Optimization
Figure 1: Left: Architecture of GOAL for single-type problems; the green components are task-specific, while the others belong to the backbone. The codebook is a shared $L{\times}D$ (resp. $\bar{L}{\times}\bar{D}$) matrix applied to the node (resp. edge) representations produced by the input adapters. Right: Architecture of GOAL for multi-type problems (with here 2 node types T1,T2); “self” and “cross” are identical multi-head mixed attention blocks (same parameters per layer for all types and tasks); “FF” for the two types are also identical Feed Forward blocks per layer. For the edges, all the type combinations are shown (here T11,T12,T21,T22) but each task can define its own. The same set of parameters is used for the two architectures.
@@ 3.1 Constructive CO problem solving
We assume that a CO problem instance is given by a pair of a finite, non-empty set $X$ of feasible solutions, and a real-valued objective function $f$ whose domain contains $X$. Solving the instance consists in finding an optimal solution, i.e. an element of $\arg\min_{x\in X}f(x)$. Constructive approaches assume that $f$ is defined not only over $X$ but over a set of partial solutions, which includes $X$ and is amenable to a step-wise construction procedure where each partial solution can be obtained by a sequence of construction steps. Thus, solving the problem instance can be formulated as a sequential decision procedure which iteratively selects the construction steps so as to reach not only a feasible but an optimal solution. That in turn can be modeled as a Markov Decision Problem (MDP) associated with the problem instance. In this paper, we use the BQ-MDPs introduced by [drakulic_bq-nco_2023] which are defined for the widespread class of tail-recursive problems  [bertsekas_dynamic_vol1_2012]. In BQ-MDPs, the actions are the construction steps, typically the selection of a node, while the states are instances.
@@ 3.2 Multi-Task CO Problems
Now, given a class $\mathcal{T}$ of tail-recursive CO problems called tasks, let $\boldsymbol{\Omega}_{t}$ for each task $t{\in}\mathcal{T}$ denote the parametric space used to represent its instances. Consider the new (multi-task) CO problem whose instances are represented in the disjoint union $\biguplus_{t\in\mathcal{T}}\boldsymbol{\Omega}_{t}$ consisting of pairs $(t,\omega)$ of a task identifier $t{\in}\mathcal{T}$ and a point $\omega{\in}\boldsymbol{\Omega}_{t}$ in its corresponding parameter space. It is easy to show that the multi-task problem is tail-recursive whenever each of its components is. And its BQ-MDP is then simply the disjoint union of the BQ-MDPs of its components.
In this paper, we study alternatives for the multi-task policy space, beyond the (trivial) Cartesian product of the policy spaces of the individual tasks, where no parameter is shared between the different tasks. The underlying assumption is that if the tasks are sufficiently similar, learning one task may elicit implicit skills useful to other tasks, which are then leveraged by parameter sharing. In the extreme, some tasks may require little training data and/or just a few shot fine-tuning to obtain acceptable performance, leveraging their similarity to more thoroughly trained tasks. Since each task has its own parametric space used to represent its instances, some part of any multi-task policy has to be task specific, and trained by processing at least a few instances of each task: pure zero-shot generalization is not possible. The situation is close to multi-lingual Natural Language Processing, where the vocabulary of each language needs to be seen at least once at training or fine-tuning. And we borrow the solution used in that context, namely adapters [bapna_simple_2019].